soybean meal is 18% protein; cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360lbs mixture that is 17% protein?

How many pounds of the cornmeal should be in the mixture?


How many pounds of the soymeal should be in the mixture?

Answer by checkley71(8403)

.18x+.09(360-x)=.17*360
.18x+32.4-.09x=61.2
.18x-.09x=61.2-32.4
.09x=28.8
x=28.8/.09
x=320 pounds of soybean meal is needed.
360-320=40 pounds of cornmeal is needed.
proof
.18*320+.09*40=61.2
57.6+3.6=61.2
61.2=61.2